# Approximations for the number of maxima and near-maxima in independent data

### v2

@@ 1.2 The absolutely continuous case
Let $X\sim\text{U}(0,1)$ have a uniform distribution on the interval $(0,1)$, so that $f(x)=1$ and $F(x)=x$ for $0\leq x\leq 1$. The uniform distribution is in the maximum domain of attraction of a Weibull (not a Gumbel) distribution, so although we do not expect a negative binomial limit for $K_{n}(a,\ell)-1$ in the case of fixed $a$ here, we may nevertheless apply our theorem to give explicit upper bounds in the approximation of $K_{n}(a,\ell)-1$ which may be useful both numerically and in determining conditions on $a=a(n)$ and $\ell=\ell(n)$ under which we may see a distribution which is close to a negative binomial. We assume that $a=a(n)\in(0,1)$ for each $n$, and for simplicity we will here assume that $\ell$ is fixed (i.e., independent of $n$). With our choice of $X$ we have that
@@
$M_{j}=n\binom{n-1}{\ell-1}\left[\int_{0}^{a}x^{n-\ell}\left(1-x\right)^{\ell-1}\,\text{d}x+a^{j}\int_{a}^{1}x^{n-\ell-j}\left(1-x\right)^{\ell-1}\,\text{d}x\right]\,.$
Since $\int_{0}^{a}x^{n-\ell}\left(1-x\right)^{\ell-1}\,\text{d}x$ is of order $O(a^{n-\ell+1}n^{-\ell})$ and $\int_{a}^{1}x^{n-\ell-1}\left(1-x\right)^{\ell-1}\,\text{d}x$ is of order $O(n^{-\ell})$, we have that $M_{1}$ is of order $O(a)$. Similarly, since $\int_{a}^{1}x^{n-\ell-2}\left(1-x\right)^{\ell-1}\,\text{d}x$ is also of order $O(n^{-\ell})$, we have that $M_{2}$ is of order $O(a^{2})$.
Recalling that $\mathbb{E}[K_{n}(a,\ell)]-1=(n-\ell)M_{1}$, if we choose $a$ such that $an\to 0$ as $n\to\infty$, then
$\beta=\frac{(n-\ell)M_{1}}{(n-\ell)M_{1}+\ell}\to 0$
and $(n-\ell-1)\frac{M_{2}}{M_{1}}-(n-\ell-2)M_{1}\to 0$. Writing $(1-\beta)^{\ell}\geq 0$, Theorem 5 then gives us the upper bound
$d_{\text{TV}}(K_{n}(a,\ell)-1,Z)\leq\frac{(n-\ell)M_{1}+\ell}{\ell}\left(\beta+(1-\beta)\left[(n-\ell-1)\frac{M_{2}}{M_{1}}-(n-\ell-2)M_{1}\right]\right)\to 0$
as $n\to\infty$ for this choice of $a=a(n)$.

### v3

@@ 1.2 The absolutely continuous case
Let $X\sim\text{U}(0,1)$ have a uniform distribution on the interval $(0,1)$, so that $f(x)=1$ and $F(x)=x$ for $0\leq x\leq 1$. The uniform distribution is in the maximum domain of attraction of a Weibull (not a Gumbel) distribution, so although we do not expect a negative binomial limit for $K_{n}(a,\ell)-1$ in the case of fixed $a$ here, we may nevertheless apply our theorem to give explicit upper bounds in the approximation of $K_{n}(a,\ell)-1$ which may be useful both numerically and in determining conditions on $a=a(n)$ and $\ell=\ell(n)$ under which we may see a distribution which is close to a negative binomial. We assume that $a=a(n)\in(0,1)$ for each $n$, and to keep the exposition concise we will here again assume that $\ell=1$; similar calculations may be carried out for other values of $\ell$. With our choice of $X$ we have that
@@
$M_{j}=n\left(\int_{0}^{a}x^{n-1}\,\text{d}x+a^{j}\int_{a}^{1}x^{n-j-1}\,\text{d}x\right)\,.$
Hence, $M_{1}=\frac{an-a^{n}}{n-1}$ and $M_{2}=\frac{a^{2}n-2a^{n}}{n-2}$. Recall that $\mathbb{E}[K_{n}(a,1)]-1=(n-1)M_{1}$. If we choose $a$ such that $an\to 0$ as $n\to\infty$, then
$\beta=\frac{(n-1)M_{1}}{(n-1)M_{1}+1}=\frac{an-a^{n}}{an-a^{n}+1}\to 0\,$,
as $n\to\infty$. From Theorem 5 we obtain the upper bound
$d_{\text{TV}}(K_{n}(a,1)-1,Z)$ $\leq(n-1)M_{1}\left(\beta+(1-\beta)\left[(n-2)\frac{M_{2}}{M_{1}}-(n-3)M_{1}\right]\right)$
$=\frac{an-a^{n}}{an-a^{n}+1}\left(\frac{(n-1)(a^{2}n-2a^{n})}{an-a^{n}}+\frac{2(an-a^{n})}{n-1}\right)\,$,
where $Z\sim\text{NB}(1,1-\beta)$. This upper bound converges to zero as $n\to\infty$ under our assumption that $an\to 0$.